OpenStudy (anonymous):

trying to simplify this square root -4/645

OpenStudy (anonymous):

\[\huge \sqrt[-4]{645}\] ?? this

OpenStudy (anonymous):



OpenStudy (anonymous):

yes i didn't know how to make the square root symbol.thanks

OpenStudy (anonymous):

In the beginning you've typed 645 now you've typed 625 ?? O_O

OpenStudy (anonymous):

which one? :)

OpenStudy (anonymous):

its 625 sorry

OpenStudy (lgbasallote):

\[-^4\sqrt{(5)(5)(5)(5)}\] factor out of the radical all the numbers that appeared 4 times

OpenStudy (anonymous):

\[\LARGE -\sqrt[4]{625}=-\sqrt[4]{5\cdot 5\cdot 5\cdot 5}\] \[\LARGE -\sqrt[4]{5^4}=...\]

OpenStudy (lgbasallote):

@Kreshnik what is the latex code for the index? i just cheated mine =)) yours seem much more professional

OpenStudy (anonymous):

thanks to both of you.

OpenStudy (anonymous):

thanks to both of you.

OpenStudy (anonymous):

\[\huge \sqrt [ x ] { y }\] appears: \[\huge \sqrt [ x ] { y }\]

OpenStudy (anonymous):

ahh.. :@

OpenStudy (lgbasallote):

lol dont put the \[ thingies hahah

OpenStudy (anonymous):

\huge \sqrt [ x ] { y } appears: \[\huge \sqrt [ x ] { y }\] :) sorry about that :)

OpenStudy (lgbasallote):

ahh thanks

OpenStudy (anonymous):

You're welcome